### v1

@@ 3.1 Basic Properties
Using the definition of $g(\mathbf{x},\mathbf{u})$ and the fact that $\mathbf{u}$ is a unit vector we immediately obtain $\langle\operatorname{grad}f(\mathbf{x}),g(\mathbf{x},\mathbf{u})\rangle=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle^{2}=\|g(\mathbf{x},\mathbf{u})\|^{2}.$
Note that if $\mathbf{u}$ is uniformly distributed on the sphere, then
@@
The image of the function $u\mapsto g(\mathbf{x},\mathbf{u})$ is a hypersphere in $\mathsf{T}_{\mathbf{x}}M$ with center $\tfrac{1}{2}\operatorname{grad}f(\mathbf{x})$ and containing the origin. Let $\mathbf{x}\in M$ be fixed and consider the map $g_{\mathbf{x}}:\mathsf{S}_{\mathbf{x}}M\to\mathsf{T}_{\mathbf{x}}M$ given by $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$. To simplify formulas, we choose an orthonormal basis $(e_{1},\ldots,e_{N})$ in $\mathsf{T}_{\mathbf{x}}M$ such that $\operatorname{grad}f(\mathbf{x})=\|\operatorname{grad}f(\mathbf{x})\|e_{N}$. In these coordinates the map $g_{\mathbf{x}}$ is given by
@@
and let $v$ be distributed on the unit sphere according to a probability measure $\mathcal{V}(\mathsf{S}_{x}M)$ such that $g$ and $\tilde{g}$ induce the same probability measure. The measure $\mathcal{V}$ is not uniform on the sphere, but it is still invariant under rotations preserving $\operatorname{grad}f({\mathbf{x}})$. We see that $v_{N}\sim 2u_{N}^{2}-1$.
Note that since the standard deviation of $u_{N}^{2}$ is larger than the expected value, Chebyshev’s inequality cannot be applied to obtain useful concentration bounds. We will use a different method in Lemma A.2.
@@
If $\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}]=\mathds{1}/N$, which is satisfied for the Haar measure, then it holds that
@@
This is a simple computation: $\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}\operatorname{grad}f(\mathbf{x})]=\frac{1}{N}\operatorname{grad}f(\mathbf{x})$. ∎
The corollary shows that in this case the projected gradient is of size $\operatorname{grad}f(\mathbf{x})/N$, and hence it is rather small for large $N$. Intuitively this happens because in high dimensions, a uniformly random unit vector $u$ will be close to orthogonal to the gradient with high probability — after all, there are $N-1$ dimensions which are orthogonal to the gradient.
For large dimension $N$ there exist good approximations of these distributions. In fact, the distribution of a coordinate of a uniformly random unit vector approximately follows a normal distribution, and hence it’s square approximates a $\chi^{2}_{1}$ distribution. See Lemma A.2 for a precise result.
Now that we better understand the iteration rule, we want to understand by what amount the objective function value is likely to decrease after a certain number of iterations. It will be useful to use normal coordinate charts at the current point $\mathbf{x}_{i}$. These charts satisfy that the metric at the origin is trivial, and that geodesics passing through the origin are straight and uniformly parametrized, see for instance [35, Prop. 5.24]. In particular, when using a normal coordinate chart about $\mathbf{x}_{i}$, the random variable $\mathbf{x}_{i+1}$, conditioned on $\mathbf{x}_{i}$, is still distributed on a hypersphere.
@@
Let $(M,g)$ be compact and let $f:M\to\mathbb{R}$ be $\ell$-smooth at $\mathbf{x}_{i}$ and with injectivity radius $\operatorname{inj}_{M}(\mathbf{x}_{i})$. Then for $\eta\leq\min(1/\ell,\operatorname{inj}_{M}(\mathbf{x}_{i}))$ it holds that
@@
In particular the function value cannot increase.
@@
We choose a normal coordinate chart about $\mathbf{x}_{i}$ and denote the coordinates by $\mathbf{\tilde{x}}$ and the function in coordinates by $\tilde{f}$. In particular $\mathbf{\tilde{x}}_{i}=0$. We may and do assume that $\nabla\tilde{f}$ is $\ell$-Lipschitz. Note that at the origin $\operatorname{grad}\tilde{f}(\mathbf{x}_{i})=\nabla\tilde{f}(\mathbf{\tilde{x}}_{i})$ and the coordinate representation of $\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i}))$ is given by $\mathbf{\tilde{x}}_{i+1}=-\eta g(\mathbf{x}_{i},\mathbf{u}_{i})$. Then by [46, Lemma 1.2.3] and Lemma 3.1 it holds that
@@
$\leq\langle\nabla\tilde{f}(\mathbf{\tilde{x}}_{1}),\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\rangle+\frac{\ell}{2}\|\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\|^{2}$
@@
This result shows that, under the assumptions of Corollary 3.3, we obtain
$\mathbb{E}[f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})]\leq-\frac{\eta}{N}\Big{(}1-\frac{\ell\eta}{2}\Big{)}\|\operatorname{grad}f(\mathbf{x}_{i})\|^{2}.$
More generally the previous result shows that the only way for the algorithm to stop improving is for the gradient to vanish: $\operatorname{grad}f(\mathbf{x}_{i})\to 0$.
@@
If $f(\mathbf{x}_{i})\to c$ then $c$ is a critical value and $\mathbf{x}_{i}$ converges almost surely to the critical set of $c$.
@@
We will show that $\Pr(f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0\wedge\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0)=0$, which immediately implies that if the values $f(\mathbf{x}_{i})$ converge, then the gradients $\operatorname{grad}f(\mathbf{x}_{i})$ converge to 00 almost surely. Note that if $\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0$, then there is some $\varepsilon>0$ such that $\|\operatorname{grad}f(\mathbf{x}_{i})\|>\varepsilon$ on some infinite subsequence. Consider the case of the Haar measure. By the above, if $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0$, then $u_{N}^{(i)}\to 0$ on the same subsequence. Since the $u_{N}^{(i)}$ are i.i.d (see Lemma A.1) and independent of the $\mathbf{x}_{i}$ the probability that this happens is 00, and this concludes the proof. The case of discrete probability distributions follows immediately form Lemma A.3. ∎
@@ 3.2 Escaping Saddle Points
The main difficulty in proving almost sure convergence of the randomly projected gradient descent algorithm is to show that it does not get stuck in a saddle point.
Recall from Lemma 3.4 that the function value cannot increase. Hence, once we reach the critical value corresponding to some saddle point, without sitting in the saddle point itself, we say that we have “passed” the saddle point as now it is impossible to converge to the saddle point in question.
@@
$f(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{n-p-q})=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}),$
with $1\leq p,q$ and $p+q\leq n$ as well as $a_{i},b_{j}>0$. Then $f$ is $\ell$-smooth with $\ell=2\max_{i,j}(a_{i},b_{j})$. Moreover let $0<\eta\leq\tfrac{1}{\ell}$ and
$\theta=\arctan\Big{(}\frac{b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}}{a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}}\Big{)}\in\big{[}0,\tfrac{\pi}{2}\big{]}.$ (8)
Then there exist constants $\varepsilon,\delta>0$ such that for all  [^4] The value $\tfrac{\pi}{3}$ is arbitrary and could be replaced by any other value in $(0,\frac{\pi}{2})$ without changing the proof. $\theta\in[0,\tfrac{\pi}{3}]$ it holds that
@@
and the value of $\ell$ follows immediately. A key property of this simplified setting is that the gradient is a linear vector field on $\mathbb{R}^{n}$, and so the entire situation is invariant under scaling. From the definition of $\theta$ it is clear that $f$ vanishes if and only if $\theta=\tfrac{\pi}{4}$.
First note that due to the scaling invariance of the algorithm and of (8), we may focus on an initial point $\mathbf{x}_{i}$ on the unit sphere $S^{n-1}$. After executing one step, we obtain a random improvement of our angle, namely $\theta_{i+1}-\theta_{i}$ which depends on $\mathbf{u}_{i}$ appearing in the update rule. Considering this improvement as a function of $\mathbf{x}_{i}$ and $\mathbf{u}_{i}$, i.e.,
$h(\mathbf{x}_{i},\mathbf{u}_{i}):=\theta_{i}-\theta_{i},$
@@
is is clear that it is a continuous function. The optimal improvement achievable for a given $\mathbf{x}_{i}$ is found by maximizing over $\mathbf{u}_{i}$. One can show that the resulting function is still continuous in $\mathbf{x}_{i}$. Moreover we claim that it is strictly positive on the sphere. Indeed, if $\theta_{i}>0$, then this follows from the choice of $\eta$. If we perform a single (deterministic) negative gradient step with step size $\eta$, none of the coordinates will change sign, and whenever $\theta\in(0,\tfrac{\pi}{2})$ its value will strictly increase. If $\theta_{i}=0$, it is clear that the value of $\theta$ will almost surely strictly increase, as the set of points $\mathbf{x}_{i+1}$ with $\theta_{i+1}=0$ form a strict subspace of $\mathbb{R}^{n}$. Taken together this shows that there is some $\varepsilon>0$ such that the optimal improvement is at least $2\varepsilon$ for every initial $\mathbf{x}_{i+1}$.
@@
It follows that for any given initial state $\mathbf{x}_{i+1}$, the probability that $\theta_{i+1}-\theta_{i}\geq\varepsilon$ is strictly positive. Indeed, $\mathbf{x}_{i+1}$ takes values on a sphere, and $\theta(\cdot)$ is a continuous function on this sphere. The preimage of $(\varepsilon,+\infty)$ on the sphere is a non-empty open set and hence has a strictly positive probability. Moreover one can show that this probability depends continuously on the initial state $\mathbf{x}_{i}$, and again we obtain a positive minimum $\delta$ on the sphere by compactness. This proves that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$ independently of the initial state.
Now consider the discrete case. Again we need to prove that there exist constants $\varepsilon,\delta>0$ such that for all $\theta\in[0,\tfrac{\pi}{3}]$ it holds that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$.
@@
The only cases where $h(\mathbf{x},\mathbf{u})=0$ (note that it cannot be negative) with probability one is when all $v_{j}(\mathbf{x})$ for $j=1,\ldots,k$ are either orthogonal to $\operatorname{grad}(\mathbf{x})$ or collinear to $\mathbf{x}$. By assumption, this cannot happen. Hence $\max_{j}h(\mathbf{x},\mathbf{v}_{j})$ is non-negative and continuous, and on the compact set $\{\mathbf{x}\in S^{n-1}:\theta(\mathbf{x})\leq\pi/3\}$ it is even strictly positive. If we denote the minimum by $\varepsilon$ and $\delta=\frac{1}{k}$, then the result follows. ∎
@@
It follows that for any $m\in\mathbb{N}$
@@
It follows immediately from Lemma 3.6 that in $N$ steps we achieve $\Pr(\theta_{i+N}-\theta_{i}\geq N\varepsilon)\geq\delta^{N}$, or put differently, setting $N=\lceil\tfrac{\pi}{3\varepsilon}\rceil$, we get
@@
In order to generalize Corollary 3.7 to the desired setting we will treat the effects of the Riemannian metric and the higher order terms of $f$ as a perturbation on the algorithm. We start with the following general perturbation result.
@@
Let $G$ be a linear vector field on $\mathbb{R}^{n}$ and let $\eta>0$ be given. Moreover let $P$ be a (not necessarily linear) vector field on $\mathbb{R}^{n}$ satisfying $\|P({\mathbf{x}})\|\leq C\|{\mathbf{x}}\|^{2}$ for some $C>0$. Then it holds that
@@
where $\|\mathds{1}+\eta G\|$ denotes the operator norm. Hence after $n$ steps, if the states remain in an $R$-ball about the origin,
$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
@@
$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|$ $\leq\|\mathds{1}+\eta G\|^{n}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+(\|\mathds{1}+\eta G\|^{n-1}+\cdots+1)CR^{2}$
$\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
@@
In our case, we can always choose coordinates in which the problem locally looks like the setting of Lemma 3.6 with a perturbation as in Lemma 3.8.
@@
Let $(M,g)$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$. If $\mathbf{x}\in M$ is a critical point of $f$, then there exists a chart $\sigma:U\to\mathbb{R}^{n}$ about $\mathbf{x}$ such that, with $\mathbf{x}=(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{n-p-q})$
$\sigma(\mathbf{x})$ $=0,$
$f\circ\sigma^{-1}(\mathbf{x})$ $=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2})+\mathcal{O}(x^{2}),$
$g_{ij}(\mathbf{x})$ $=\delta_{ij}+\mathcal{O}(x),$
@@
Choose coordinates about $\mathbf{x}$ such that the metric is Euclidean at the $x$ (e.g. normal coordinates), then use orthogonal transformation to diagonalize the Hessian of $f$ at $x$. ∎

### v2

@@ 3.1 Basic Properties
Using the definition of $g(\mathbf{x},\mathbf{u})$ and the fact that $\mathbf{u}$ is a unit vector, we immediately obtain $\langle\operatorname{grad}f(\mathbf{x}),g(\mathbf{x},\mathbf{u})\rangle=\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle^{2}=\|g(\mathbf{x},\mathbf{u})\|^{2}.$
Corollary 3.4.
Note that if $\mathbf{u}$ is uniformly distributed (w.r.t. the Haar measure) on the sphere, then
@@
The image of the function $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$ is a hypersphere in $\mathsf{T}_{\mathbf{x}}M$ with center $\tfrac{1}{2}\operatorname{grad}f(\mathbf{x})$ and intersecting the origin: Indeed, let $\mathbf{x}\in M$ be fixed and consider the map $g_{\mathbf{x}}:\mathsf{S}_{\mathbf{x}}M\to\mathsf{T}_{\mathbf{x}}M$ given by $\mathbf{u}\mapsto g(\mathbf{x},\mathbf{u})$. To simplify formulas, we choose an orthonormal basis $(e_{1},\ldots,e_{N})$ in $\mathsf{T}_{\mathbf{x}}M$ such that $\operatorname{grad}f(\mathbf{x})=\|\operatorname{grad}f(\mathbf{x})\|e_{N}$. In such coordinates the map $g_{\mathbf{x}}$ is given by
@@
and let $\mathbf{v}$ be distributed on the unit sphere according to a probability measure $\mathcal{V}(\mathsf{S}_{x}M)$ such that $g$ and $\tilde{g}$ induce the same probability measure. The measure $\mathcal{V}$ is not uniform on the sphere, but it is still invariant under rotations preserving $\operatorname{grad}f({\mathbf{x}})$. We see that $v_{N}\sim 2u_{N}^{2}-1$. More details on the distribution of $u_{N}$ and $u_{N}^{2}$ can be found in Lems. A.1 and A.2.
Note that since the standard deviation of $u_{N}^{2}$ is larger than the expected value, Chebyshev’s inequality cannot be applied to obtain useful concentration bounds. For large dimension $N$, there exist good approximations of these distributions. In fact, the distribution of a coordinate of a uniformly random unit vector approximately follows a normal distribution, and hence it’s square approximates a $\chi^{2}_{1}$ distribution. See Lem. A.2 for a precise result.
@@
If $\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}]=\mathrm{I}_{N}/N$, which is satisfied for the Haar measure, then it holds that
@@
We compute: $\mathbb{E}[g(\mathbf{x},\mathbf{u})]=\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}\operatorname{grad}f(\mathbf{x})]=\frac{1}{N}\operatorname{grad}f(\mathbf{x})$. ∎
The corollary shows that in this case the projected gradient is of size $\operatorname{grad}f(\mathbf{x})/N$, and hence it is rather small for large $N$. Intuitively, this happens because in high dimensions, a uniformly random unit vector $u$ will be close to orthogonal to the gradient with high probability — after all, there are $N-1$ dimensions which are orthogonal to the gradient.
Now that we better understand the iteration rule, we want to understand by what amount the objective function value is likely to decrease after a certain number of iterations. It will be useful to use normal coordinate charts at the current point $\mathbf{x}_{i}$. These charts satisfy that the metric at the origin is trivial, and that geodesics passing through the origin are straight and uniformly parametrized, see for instance [42, Prop. 5.24]. In particular, when using a normal coordinate chart about $\mathbf{x}_{i}$, the random variable $\mathbf{x}_{i+1}$, conditioned on $\mathbf{x}_{i}$, is still distributed on a hypersphere.
@@
Let $(M,\mathsf{g})$ be compact with injectivity radius $r_{M}$ and let $f:M\to\mathbb{R}$ be $\ell$-smooth. Then for $\eta\leq\min(1/\ell,r_{M})$ it holds that
@@
In particular the function value is surely non-increasing.
@@
We choose a normal coordinate chart about $\mathbf{x}_{i}$ and denote the coordinates by $\mathbf{\tilde{x}}$ and the function in coordinates by $\tilde{f}$. In particular $\mathbf{\tilde{x}}_{i}=0$. Note that at the origin $\operatorname{grad}\tilde{f}(\tilde{\mathbf{x}}_{i})=\nabla\tilde{f}(\mathbf{\tilde{x}}_{i})$ and the coordinate representation of $\mathbf{x}_{i+1}=\exp_{\mathbf{x}_{i}}(-\eta g(\mathbf{x}_{i},\mathbf{u}_{i}))$ is given by $\mathbf{\tilde{x}}_{i+1}=-\eta\tilde{g}(\mathbf{x}_{i},\mathbf{u}_{i})$, where $\tilde{g}(\mathbf{x}_{i},\mathbf{u}_{i})$ is the chart representation of $g(\mathbf{x}_{i},\mathbf{u}_{i})$. Then, by the proof of [55, Lemma 1.2.3] and Lem. 3.3, it holds that
@@
$\leq\langle\nabla\tilde{f}(\mathbf{\tilde{x}}_{i}),\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\rangle+\frac{\ell}{2}\|\mathbf{\tilde{x}}_{i+1}-\mathbf{\tilde{x}}_{i}\|^{2}$
@@
This result shows that, under the assumptions of Cor. 3.6, we obtain the conditional expectation
$\mathbb{E}[f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\,|\,\mathbf{x}_{i}]\leq-\frac{\eta}{N}\Big{(}1-\frac{\ell\eta}{2}\Big{)}\|\operatorname{grad}f(\mathbf{x}_{i})\|^{2}.$
More generally the previous result shows that the only way for the algorithm to stop improving is for the gradient to vanish: $\operatorname{grad}f(\mathbf{x}_{i})\to 0$ as $i\to\infty$.
@@
Under the assumptions of Lem. 3.7, the function values converge (surely) to some value $c$. Moreover, the following assertions hold almost surely
  1. (i)
    $c$ is a critical value of $f$.
  2. (ii)
    $\mathbf{x}_{i}$ converges to the critical set of $c$, i.e. to the set of critical point with $f(\mathbf{x})=c$.
@@
By Lem. 3.7 the function values are non-increasing and bounded below, hence they converge to some value $c$. We will show that $\Pr(\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0)=0$, which immediately implies the result. Since $f(\mathbf{x}_{i+1})-f(\mathbf{x}_{i})\to 0$, Lem. 3.7 shows that $\langle\operatorname{grad}f(\mathbf{x}_{i}),g(\mathbf{x}_{i},\mathbf{u}_{i})\rangle\to 0$.
Note that if $\operatorname{grad}f(\mathbf{x}_{i})\nrightarrow 0$, then there is some $m\in\mathbb{N}$ such that $\|\operatorname{grad}f(\mathbf{x}_{i})\|>\tfrac{1}{m}$ on some infinite subsequence. Let $z_{j}^{m}=\langle\mathbf{u}_{i(j)},\operatorname{grad}f(\mathbf{x}_{i(j)})\rangle/\|\operatorname{grad}f(\mathbf{x}_{i(j)})\|$ where $i(j)$ is the $j$-th index such that $\|\operatorname{grad}f(\mathbf{x}_{i(j)})\|\geq\tfrac{1}{m}$. Note that
$\langle\operatorname{grad}f(\mathbf{x}_{i(j)}),g(\mathbf{x}_{i(j)},\mathbf{u}_{i(j)})\rangle=(z_{j}^{m}\|\operatorname{grad}f(\mathbf{x}_{i(j)})\|)^{2}\,.$
Hence, if the gradients do not converge to 00, then, for some $m$, it must hold that $|z_{j}^{m}|\to 0$.
We claim that there exist $10<\varepsilon,\delta<$ 1 such that for all $\mathbf{x}\in M$ with $f(\mathbf{x})\leq f(\mathbf{x}_{0})$ it holds that
$\Pr(|\langle\mathbf{u},\operatorname{grad}f(\mathbf{x})\rangle|\leq\varepsilon\|\operatorname{grad}f(\mathbf{x})\|)\leq\delta\,.$
If $\mathbf{u}\sim\mathcal{U}(\mathsf{S}_{\mathbf{x}}M)$ this follows from Lem. A.1, if $\mathbf{u}\sim\mathcal{D}(\mathsf{S}_{\mathbf{x}}M)$ this follows from Lem. 3.2.
Given $m,n,k\in\mathbb{N}$ with $1/n<\varepsilon$, and using the chain rule, Markovianity, and the claim above, we compute that
$\Pr(\forall j>k:|z_{j}^{m}|\leq\tfrac{1}{n})=\prod_{j=k}^{\infty}\Pr(|z_{j}^{m}|\leq\tfrac{1}{n}\;\big{|}\;\forall l\in\{j,\ldots,k\}:|z_{l}^{m}|\leq\tfrac{1}{n})=0\,.$
For any fixed value of $n$, we can use a union bound over all $m,k\geq 1$ and obtain
$\Pr(\exists m,k\,\forall j>k:|z_{j}^{m}|\leq\tfrac{1}{n})=0\,,$
which implies that
$\Pr(\exists m:|z_{j}^{m}|\to 0\text{ as }j\to\infty)=0\,,$
as desired. ∎
@@ 3.2 Escaping Saddle Points
The main difficulty in proving almost sure convergence to a local minimum of the randomly projected gradient descent algorithm is to show that it does not get stuck in a saddle point.
Recall from Lem. 3.7 that the function value cannot increase. Hence, once we cross the critical value corresponding to some saddle point, we say that we have passed the saddle point, as now it is impossible to converge to the saddle point in question.
@@
$f(x,y_{,}z)=a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}),$
with $x\in\mathbb{R}^{p}$, $y\in\mathbb{R}^{q}$, $z\in\mathbb{R}^{N-p-q}$, and $p,q\geq 1$ as well as $a_{i},b_{j}>0$. Then $f$ is $\ell$-smooth with $\ell=2\max\{a_{1},\dots,a_{p},b_{1},\dots,b_{q}\}$. Moreover, let $0<\eta\leq\tfrac{1}{\ell}$ and define
$\theta:=\arctan\Big{(}\frac{b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2}}{a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}}\Big{)}\in\big{[}0,\tfrac{\pi}{2}\big{]}\quad\text{for}\quad(x_{1},\dots,x_{p},y_{1},\dots,y_{q})\neq 0.$ (8)
Then there exist constants $\varepsilon,\delta>0$ such that for all  [^10] The value $\tfrac{\pi}{3}$ is arbitrary and could be replaced by any other value in $(\tfrac{\pi}{4},\frac{\pi}{2})$ without changing the proof. $\theta\in[0,\tfrac{\pi}{3}]$ it holds that
@@
and the value of $\ell$ follows immediately. A key property of this simplified setting is that the gradient is a linear vector field on $\mathbb{R}^{N}$, and so the entire situation is invariant under scaling. Moreover, from the definition of $\theta$ it is clear that $f$ vanishes if and only if $\theta=\tfrac{\pi}{4}$ or $(x_{1},\dots,x_{p},y_{1},\dots,y_{q})=0$. Due to the scaling invariance of the algorithm and of (8), we may focus on an initial point $\mathbf{x}_{i}=(x_{i},y_{i},z_{i})$ on the unit sphere $S^{N-1}$. After executing one step, we obtain a random improvement of our angle, namely $\theta_{i+1}-\theta_{i}$ which depends on $\mathbf{u}_{i}$ appearing in the update rule.
First let us consider this improvement in the uniform Case I as a function of $\mathbf{x}_{i}$ and $\mathbf{u}_{i}$, i.e.,
$h(\mathbf{x}_{i},\mathbf{u}_{i}):=\theta_{i+1}-\theta_{i}\,.$
@@
It is clear that this is a continuous function. The optimal improvement achievable for a given $\mathbf{x}_{i}$ is found by maximizing over $\mathbf{u}_{i}$. One can show that the resulting function is still continuous in $\mathbf{x}_{i}$. Moreover we claim that it is strictly positive on the compact set $K_{\pi/3}:=\{\mathbf{x}\in S^{n-1}:\theta(\mathbf{x})\leq\pi/3\}$. Indeed, if $\theta_{i}>0$, then this follows from the choice of $\eta$. If we perform a single (deterministic) negative gradient step with step size $\eta$, none of the coordinates will change sign, and whenever $\theta_{i}\in(0,\tfrac{\pi}{3}]$ its value will strictly increase. If $\theta_{i}=0$, it is clear that the value of $\theta_{i}$ will almost surely strictly increase, as the set of points $\mathbf{x}_{i+1}$ with $\theta_{i+1}=0$ form a strict subspace of $\mathbb{R}^{n}$. Taken together this shows that there is some $\varepsilon>0$ such that the optimal improvement is at least $2\varepsilon$ for every initial $\mathbf{x}_{i}\in K_{\pi/3}$.
@@
It follows that for any given initial state $\mathbf{x}_{i}\in K_{\pi/3}$, the probability that $\theta_{i+1}-\theta_{i}\geq\varepsilon$ is strictly positive. Indeed, $h(\mathbf{x}_{i},\cdot)$ is a continuous function on this sphere and its preimage of $(\varepsilon,+\infty)$ on the sphere is a non-empty open set and hence has a strictly positive probability (with respect to the Haar measure). Moreover one can show that this probability depends continuously on the initial state $\mathbf{x}_{i}$, and again we obtain a positive minimum $\delta$ on the $K_{\theta}$ by compactness. This proves that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$ independently of the initial state $\mathbf{x}\in K_{\pi/3}$.
@@
Next consider the discrete Case II. Again we need to prove that there exist constants $\varepsilon,\delta>0$ such that for all $\theta\in[0,\tfrac{\pi}{3}]$ it holds that $\Pr(\theta_{i+1}-\theta_{i}\geq\varepsilon)\geq\delta$. The only cases where $h(\mathbf{x},\mathbf{u})=0$ (note that it cannot be negative) with probability one is when all $v_{j}(\mathbf{x})$ for $j=1,\ldots,k$ are either orthogonal to $\operatorname{grad}(\mathbf{x})$ or collinear to $\mathbf{x}$. By Assumption 3.1, this cannot happen. Hence $\max_{j}h(\mathbf{x},\mathbf{v}_{j})$ is non-negative and continuous, and on the compact set $K_{\pi/3}$ it is even strictly positive. This concludes the proof. ∎
As a corollary, we are able to show that after sufficiently many steps, the algorithm passes the saddle point with high probability.
@@
It follows that for any $m\in\mathbb{N}$, it holds that
@@
It follows immediately from Lem. 3.9 that in $n$ steps we achieve $\Pr(\theta_{i+n}-\theta_{i}\geq n\varepsilon)\geq\delta^{n}$, or put differently, setting $n=\lceil\tfrac{\pi}{3\varepsilon}\rceil$, we get
@@
In order to generalize Cor. 3.10 to the desired setting we will treat the effects of the Riemannian metric and the higher order terms of $f$ as a perturbation on the algorithm. We start with the following general perturbation result.
@@
Let $G$ be a linear vector field on $\mathbb{R}^{n}$ and let $\eta>0$ be given. Moreover let $P$ be any other vector field on $\mathbb{R}^{n}$ satisfying $\|P({\mathbf{x}})\|\leq C\|{\mathbf{x}}\|^{2}$ for some $C>0$. Then it holds that
@@
where $\|\mathds{1}+\eta G\|$ denotes the operator norm. Let ${\mathbf{x}}_{i+1}={\mathbf{x}}_{i}-\eta G{\mathbf{x}}_{i}$ and ${{\mathbf{y}}_{i+1}}={\mathbf{y}}_{i}-\eta(G+P){\mathbf{y}}_{i}$. Hence after $n$ steps, if the states remain in an $R$-ball about the origin, then
$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i}-{\mathbf{y}}_{i}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
@@
$\|{\mathbf{x}}_{i+n}-{\mathbf{y}}_{i+n}\|$ $\leq\|\mathds{1}+\eta G\|^{n}\|{\mathbf{x}}_{i}-{\mathbf{y}}_{i}\|+(\|\mathds{1}+\eta G\|^{n-1}+\cdots+1)CR^{2}$
$\leq\|\mathds{1}+\eta G\|^{n}\Big{(}\|{\mathbf{x}}_{i}-{\mathbf{y}}_{i}\|+\frac{CR^{2}}{\|\mathds{1}+\eta G\|-1}\Big{)}.$
@@
In our case, we can always choose coordinates in which the problem locally looks like the setting of Lem. 3.9 with a perturbation as in Lem. 3.11.
@@
Let $(M,\mathrm{g})$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be a Morse–Bott function on $M$. If $\mathbf{x}_{0}\in M$ is a critical point of $f$, then there exists a chart $\sigma:U\to\mathbb{R}^{n}$ about $\mathbf{x}_{0}$ such that, with $\tilde{\mathbf{x}}=\sigma(\mathbf{x})=(x_{1},\ldots,x_{p},y_{1},\ldots,y_{q},z_{1},\ldots,z_{N-p-q})$,
$\sigma(\mathbf{x}_{0})$ $=0,$
$f\circ\sigma^{-1}(\tilde{\mathbf{x}})$ $=c+a_{1}x_{1}^{2}+\ldots+a_{p}x_{p}^{2}-(b_{1}y_{1}^{2}+\ldots+b_{q}y_{q}^{2})+\mathcal{O}(\|\tilde{\mathbf{x}}\|^{2}),$
$\mathrm{g}_{ij}(\tilde{\mathbf{x}})$ $=\delta_{ij}+\mathcal{O}(\|\tilde{\mathbf{x}}\|),$
@@
Choose coordinates about $\mathbf{x}_{0}$ such that the metric is Euclidean at $\tilde{\mathbf{x}}_{0}$ (e.g. normal coordinates), then use an orthogonal transformation to diagonalize the Hessian of $f$ at $\tilde{\mathbf{x}}_{0}$. ∎